### v1

Abstract
The randomized singular value decomposition (SVD) has become a popular approach to computing cheap, yet accurate, low-rank approximations to matrices due to its efficiency and strong theoretical guarantees. Recent work by Boullé and Townsend (FoCM, 2023) presents an infinite-dimensional analog of the randomized SVD to approximate Hilbert–Schmidt operators. However, many applications involve computing low-rank approximations to symmetric positive semi-definite matrices. In this setting, it is well-established that the randomized Nyström approximation is usually preferred over the randomized SVD. This paper explores an infinite-dimensional analog of the Nyström approximation to compute low-rank approximations to non-negative self-adjoint trace-class operators. We present an analysis of the method and, along the way, improve the existing infinite-dimensional bounds for the randomized SVD. Our analysis yields bounds on the expected value and tail bounds for the Nyström approximation error in the operator, trace, and Hilbert-Schmidt norms. Numerical experiments for simple integral operators validate the proposed framework.
keywords:
Low-rank approximation, randomized numerical linear algebra, Nyström approximation, Hilbert-Schmidt operators
{MSCcodes}
65F30, 68W20, 65R10
1 Introduction
Randomized techniques [19, 20, 30, 39] are becoming increasingly popular for computing low-rank approximations to matrices. Most notably, the randomized singular value decomposition (SVD) [20] has evolved into one of the primary choices, due to its simplicity, performance, and reliability. In its most basic form, the randomized SVD performs the approximation $\bm{Q}\bm{Q}^{*}\bm{A}\approx\bm{A}$, where $\bm{Q}$ is an orthonormal basis for the range of $\bm{A}\bm{\Omega}$, with $\bm{\Omega}$ being a tall and skinny random sketch matrix. In many applications of low-rank approximation, such as $k$-means clustering [41], PCA [44], and Gaussian process regression [18], it is known that $\bm{A}$ is symmetric positive semi-definite (SPSD). In this case, one usually prefers the so-called randomized Nyström approximation [19]
$\widehat{\bm{A}}:=\bm{A}\bm{\Omega}(\bm{\Omega}^{*}\bm{A}\bm{\Omega})^{\dagger}\bm{\Omega}^{*}\bm{A}\approx\bm{A},$ (1)
where $\bm{\Omega}$ is, again, a random sketch matrix and $\dagger$ denotes the Moore–Penrose pseudoinverse of a matrix. This approximation has received significant attention in the literature [19, 32, 39] and, like the randomized SVD, it enjoys strong theoretical guarantees. With the same number of matrix-vector products with $\bm{A}$, the randomized Nyström approximation is provably more efficient than the randomized SVD applied to an SPSD matrix; see e.g. [38, Lem. 5.2] and [33, Thms. 3.2 and 3.5].
Recently, Boullé and Townsend [7, 8] generalized the randomized SVD from matrices to Hilbert–Schmidt operators. Subsequent works [6, 10] employed this infinite-dimensional generalization of the randomized SVD to learn Green’s functions associated with an elliptic or parabolic partial differential equations (PDE) from a few solutions of the PDE. This approach uses hierarchical low-rank techniques and exploits the fact that Green’s functions are smooth away from the diagonal and therefore admit accurate off-diagonal low-rank approximations [3, 4]. The connection between numerical linear algebra and operator learning [9] motivates the development of infinite-dimensional analog of low-rank approximation algorithms, such as the randomized Nyström approximation, to derive sample complexity estimates of neural operator methods [24, 26, 27] in scientific machine learning. Other applications, like Gaussian process regression and Support Vector Machines [16, 18, 31, 40, 42, 43], involve integral operators that feature positive and globally smooth kernels. In turn, the operator is not only self-adjoint and positive but it also allows for directly applying low-rank approximation, without the need to resort to hierarchical techniques. Given existing results on matrices, it would be sensible to use an infinite-dimensional extension of the randomized Nyström approximation in such situations.
Remark 1.1.
The use of the term Nyström approximation in some of the recent literature and this work deviates somewhat from its original meaning, where the matrix $\bm{\Omega}$ in (1) is required to consist of columns of the identity matrix $\bm{I}$ and, in turn, $\bm{A}\bm{\Omega}$ consists of selected columns of $\bm{A}$. This form of Nyström approximation frequently appears in applications involving kernel matrices; see, e.g., [2, 11, 12, 14, 43]. Following [19], we interpret the term Nyström approximation in a broader sense that allows for other choices of matrices $\bm{\Omega}$, including Gaussian random matrices.
1.1 Contributions
In this work, we present and analyze an infinite-dimensional extension of the randomized Nyström approximation for computing low-rank approximations to self-adjoint, positive, trace class operators. A difficulty in the design and analysis of our extension is that existing results of the finite-dimensional case (1) usually assume that the columns of $\bm{\Omega}$ are drawn from a standard Gaussian multivariate distribution $\mathcal{N}(\bm{0},\bm{I})$, which does not have a practically meaningful infinite-dimensional analog. Following [8], the columns of $\bm{\Omega}$ are replaced with random fields drawn from a Gaussian process $\mathcal{GP}(0,K)$, which is an infinite-dimensional analog of a non-standard multivariate Gaussian distribution.
Therefore, to analyze the infinite-dimensional generalization of the Nyström approximation, we will proceed through an analysis of the finite-dimensional Nyström approximation (1) when the columns of $\bm{\Omega}$ are drawn from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$, for some general symmetric positive semi-definite matrix $\bm{K}$. We derive expectation and probability bounds between the matrix and its approximant in the Frobenius, spectral, and nuclear norms in Theorems 2.1, 2.4, and 2.6. Then, we use continuity arguments to obtain bounds on the infinite-dimensional generalization of Eq. 1 in Theorem 3.4. As a byproduct of our analysis, we improve the analysis of the infinite-dimensional analog of the randomized SVD. In particular, unlike the bounds presented in [8], our improved bounds coincide with the bounds obtained by Halko, Martinsson, and Tropp [20] in the finite-dimensional case when $\bm{K}$ is chosen as the identity matrix.
1.2 Organization of the paper
The paper is organized as follows. In Section 2 we present the setting for the Nyström approximation in finite dimensions, give an overview of Gaussian random vectors, and provide the analysis of the Nyström approximation for matrices with correlated Gaussian sketches. We then present the generalization of the Nyström approximation to Hilbert-Schmidt operators in Section 3. Next, we conclude with numerical experiments in Section 4.
2 Nyström approximation in finite dimensions with correlated Gaussian sketches
In this section, we present an analysis of the randomized Nyström approximation defined in (1), where the columns of $\bm{\Omega}$ are independent and identically distributed (i.i.d.) random vectors with distribution $\mathcal{N}(\bm{0},\bm{K})$ for some SPSD covariance matrix $\bm{K}$. We derive bounds in the Frobenius, spectral, and nuclear norms, which we denote respectively by $\|\bm{A}\|_{\textup{F}}=(\sum_{i=1}^{n}\sigma_{i}^{2})^{1/2}$, $\|\bm{A}\|_{2}=\sigma_{1}$, and $\|\bm{A}\|_{*}=\sum_{i=1}^{n}\sigma_{i}$, where $\sigma_{1}\geq\sigma_{2}\geq\ldots,\geq\sigma_{n}$ are the singular values of $\bm{A}$. More generally, $\|\bm{A}\|_{(s)}=(\sum_{i=1}^{n}\sigma_{i}^{s})^{1/s}$ denotes the Schatten $s$-norm, which coincides with the Frobenius norm when $s=2$ and the nuclear norm when $s=1$.
We consider a symmetric positive semi-definite (SPSD) matrix $\bm{A}\in\mathbb{R}^{n\times n}$ with rank greater than $k\geq 1$ and eigenvalue decomposition $\bm{A}=\bm{U}\bm{\Sigma}\bm{U}^{*}$, partitioned as
$\bm{A}=\begin{bmatrix}\bm{U}_{1}&\bm{U}_{2}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{1}&\bm{0}\\ \bm{0}&\bm{\Sigma}_{2}\end{bmatrix}\begin{bmatrix}\bm{U}_{1}^{*}\\ \bm{U}_{2}^{*}\end{bmatrix},\quad\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k}),\quad\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\ldots,\sigma_{n}),$ (2)
where $\bm{U}_{1}\in\mathbb{R}^{n\times k}$ contains the first $k$ columns of $\bm{U}$ and $\sigma_{1}\geq\cdots\geq\sigma_{n}\geq 0$ are the eigenvalues of $\bm{A}$. For an oversampling parameter $p$ and an arbitrary sketch matrix $\bm{\Omega}\in\mathbb{R}^{n\times(k+p)}$, we denote the Nyström approximation of $\bm{A}$ defined in (1) by $\widehat{\bm{A}}$.
2.1 Distribution of Gaussian sketches
To introduce some notation and basic background material, we let
$\bm{\Omega}_{1}=\bm{U}_{1}^{*}\bm{\Omega}\in\mathbb{R}^{k\times(k+p)},\quad\bm{\Omega}_{2}=\bm{U}_{2}^{*}\bm{\Omega}\in\mathbb{R}^{(n-k)\times(k+p)},$
and, to analyze their distribution, we partition the matrix $\widetilde{\bm{K}}=\bm{U}^{*}\bm{K}\bm{U}$ as
$\widetilde{\bm{K}}=\begin{bmatrix}\bm{U}_{1}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{1}^{*}\bm{K}\bm{U}_{2}\\ \bm{U}_{2}^{*}\bm{K}\bm{U}_{1}&\bm{U}_{2}^{*}\bm{K}\bm{U}_{2}\end{bmatrix}=:\begin{bmatrix}\widetilde{\bm{K}}_{11}&\widetilde{\bm{K}}_{21}^{*}\\ \widetilde{\bm{K}}_{21}&\widetilde{\bm{K}}_{22}\end{bmatrix},\quad\widetilde{\bm{K}}_{11}\in\mathbb{R}^{k\times k}.$ (3)
We assume that $\widetilde{\bm{K}}_{11}$ is invertible, which allows us to define the Schur complement $\widetilde{\bm{K}}_{22.1}=\widetilde{\bm{K}}_{22}-\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}$. By well-known properties of Gaussian random vectors [21, Thm. 5.2], $\bm{\omega}\sim\mathcal{N}(0,\bm{K})$ implies $\bm{U}^{*}\bm{\omega}\sim\mathcal{N}(0,\widetilde{\bm{K}})$ and, thus, the marginal distributions
$\bm{\omega}_{1}:=\bm{U}_{1}^{*}\bm{\omega}\sim\mathcal{N}(0,\widetilde{\bm{K}}_{11}),\quad\bm{\omega}_{2}=\bm{U}_{2}^{*}\bm{\omega}\sim\mathcal{N}(0,\widetilde{\bm{K}}_{22}).$
In particular, the columns of $\bm{\Omega}_{1}$ are i.i.d. as $N(\bm{0},\widetilde{\bm{K}}_{11})$ or, in other words, $\bm{\Omega}_{1}=\smash{\widetilde{\bm{K}}_{11}^{1/2}\bm{X}}$ with the standard Gaussian matrix $\bm{X}\in\mathbb{R}^{k\times(k+p)}$. Because $\widetilde{\bm{K}}_{11}$ is invertible, this shows that, with probability one, $\bm{\Omega}_{1}$ has full rank and possesses a right inverse $\bm{\Omega}_{1}^{\dagger}$, which satisfies $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$. Finally, the conditional probability distribution of $\bm{\omega}_{2}$ given $\bm{\omega}_{1}=\bm{x}_{1}$ is normal with mean $\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{x}_{1}$ and covariance matrix $\widetilde{\bm{K}}_{22.1}$ [21, Thm. 5.3].
2.2 Nyström method with correlated Gaussian sketches
Similarly to the error analysis for the randomized SVD with correlated input vectors [7, 8], our error bounds for the Nyström approximation depend on prior information contained in the representation Eq. 3 of $\bm{K}$ with respect to the eigenvectors of $\bm{A}$. This is measured through the following two quantities:
$\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},$ (4)
where $\xi\in\{\textup{F},2,*\}$ such that $\|\cdot\|_{\xi}$ dictates the choice of norm. The following theorem states our main result concerning the approximation error of the Nyström approximation in the Frobenius norm. Similar results are derived in the spectral and nuclear norms but are deferred to Section 2.2.3 for clarity of exposition.
Theorem 2.1 (Frobenius norm).
Let $\bm{A}$ be an $n\times n$ SPSD matrix, $2\leq k\leq\operatorname{rank}(\bm{A})$ be a target rank, and $p\geq 4$ an oversampling parameter. Let $\bm{\Omega}$ be a random sketch matrix with the columns i.i.d. $\mathcal{N}(\bm{0},\bm{K})$, where the covariance matrix $\bm{K}$ is such that the matrix $\widetilde{\bm{K}}_{11}$ defined in Eq. 3 is invertible. Then, the Nyström approximation $\widehat{\bm{A}}$ satisfies
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}]\leq\left(1+2\delta_{k}^{(\textup{F})}+2\sqrt{c_{1}}\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2\sqrt{c_{2}}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (5)
where $c_{1}=\mathcal{O}(k^{2}/p^{2})$, $c_{2}=\mathcal{O}(k^{2}/p^{2})$ are constants defined later in Eq. 9. Let $u,t\geq 1$, then
$\|\bm{A}-\widehat{\bm{A}}\|_{\textup{F}}$ $\leq\|\bm{\Sigma}_{2}\|_{\textup{F}}+4\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+4t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}$ (6)
$\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}$
holds with probability at least $1-3t^{-p}-e^{-u^{2}/2}$. Here, $d_{1}=\mathcal{O}(k/p)$, $d_{2}=\mathcal{O}(k/p)$, and $d_{3}=\mathcal{O}(k^{3/2}/p)$ are constants defined in Eq. 12.
The proof of Theorem 2.1 is based on an existing structural bound:
$\|\bm{A}-\widehat{\bm{A}}\|\leq\|\bm{\Sigma}_{2}\|+\|(\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|,$ (7)
[…]
2.2.1 Proof of the expectation bound
To prove inequality Eq. 5 of Theorem 2.1, we use the $L_{q}$ norm of a random variable $Z$ defined as $\mathbb{E}^{q}(Z)=\mathbb{E}[|Z|^{q}]^{1/q}$. After taking expectation with respect to $\bm{\Omega}$ and applying Hölder’s inequality, the second term in the bound Eq. 7 for the Frobenius norm satisfies
$\mathbb{E}[\|(\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\bm{\Omega}}_{2}\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}]=\mathbb{E}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}]\leq\big{(}{\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]}\big{)}^{2}.$
To proceed from here, we recall that Section 2.1 provides the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ as $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix. Using the triangle inequality for the $L_{4}$ norm and $\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}=\bm{I}_{k}$ (which holds with probability $1$), we obtain that
$\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$ $=\mathbb{E}_{\bm{\Omega}_{1},\bm{\Psi}}^{4}[\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}+\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]$ (8)
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\big{(}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\mid\bm{\Omega}_{1}\big{]}\big{]}\big{)}^{1/4},$
To bound the second term, we first apply Eq. 32b,
$\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\mid\bm{\Omega}_{1}\big{]}=\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}\big{(}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}\big{)}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4},$
and then take expectation with respect to $\bm{\Omega}_{1}$ using Lemma A.5:
$\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\mid\bm{\Omega}_{1}\big{]}\big{]}\leq c_{1}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4},$
where
$c_{1}=k\frac{(p-1)(k+1)+2}{p(p-1)(p-3)},\quad\text{and}\quad c_{2}=k\frac{k+p-1}{p(p-1)(p-3)}.$ (9)
Inserting the inequality above into Eq. 8 and using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2}$ gives
$\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]\leq$
$\quad\qquad\big{(}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1/2}\|_{(4)}+\big{(}c_{1}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}+c_{2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\big{)}^{1/4}\big{)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2},$
Finally, inserting the covariance quality factors (see Eq. 4) leads to the following inequality,
$\big{(}\mathbb{E}^{4}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}]\big{)}^{2}\leq\left(2\delta_{k}^{(\textup{F})}+2\sqrt{c_{1}}\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2\sqrt{c_{2}}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$
where we used $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and the subadditivity of the square-root. This concludes the proof of°Eq. 5.
2.2.2 Proof of the tailbound
We begin the proof of the tailbound (6) of Theorem 2.1 by a concentration inequality on norms of shifted and rescaled Gaussian matrices.
Lemma 2.2.
Let $\bm{\Psi}$ be a standard Gaussian matrix and let $\bm{B},\bm{C}$, $\bm{D}$ be fixed matrices of matching sizes. Let $s\geq 2$. Then
$\mathbb{P}\left\{\|\bm{B}+\bm{C}\bm{\Psi}\bm{D}\|_{(s)}\geq\mathbb{E}\left[\|\bm{B}+\bm{C}\bm{\Psi}\bm{D}\|_{(s)}\right]+\|\bm{C}\|_{2}\|\bm{D}\|_{2}u\right\}\leq e^{-u^{2}/2}.$
holds for every $u\geq 1$.
Proof 2.3.
Given $h(\bm{X}):=\|\bm{B}+\bm{C}\bm{X}\bm{D}\|_{(s)}$, we have
$|h(\bm{X})-h(\bm{Y})|\leq\|\bm{C}(\bm{X}-\bm{Y})\bm{D}\|_{(s)}\leq\|\bm{C}\|_{2}\|\bm{D}\|_{2}\|\bm{X}-\bm{Y}\|_{(s)}\leq\|\bm{C}\|_{2}\|\bm{D}\|_{2}\|\bm{X}-\bm{Y}\|_{\textup{F}},$
where we used that the Frobenius norm is larger than any Schatten-$s$ norm for $s\geq 2$. In other words, $h$ is Lipschitz continuous with Lipschitz constant $\|\bm{C}\|_{2}\|\bm{D}\|_{2}$. This allows us to apply a concentration result for functions of Gaussian matrices [20, Prop. 10.3], which yields the statement of the lemma.
To obtain Eq. 6, it suffices to derive a tailbound for the term $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$ in the structural bound Eq. 7 (in the Frobenius norm). Using that $\bm{\Omega}_{2}|\bm{\Omega}_{1}\sim\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}+\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}$ with a standard Gaussian matrix $\bm{\Psi}$, Lemma 2.2 yields the following tailbound for $\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}$ conditioned on $\bm{\Omega}_{1}$:
$\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$ (10)
In analogy to (8), combining the triangular inequality for the Schatten-$4$ norm and Jensen’s inequality yields
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\mathbb{E}_{\bm{\Psi}}^{4}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}$
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}+\left(\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{4}\big{(}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}\big{)}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{4}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\right)^{1/4}.$
Using, again, the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and the subadditivity of the square-root, we have
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid\bm{\Omega}_{1}\big{]}^{2}$ $\leq 2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{(4)}^{2}+2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{(4)}^{2}(\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}+\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2})$ (11)
$\quad+2\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}^{2}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}.$
To control the norms of $\bm{\Omega}_{1}^{\dagger}$ in our bounds, we condition on the event $E_{t}$ that the following three inequalities are satisfied:
$\|\bm{\Omega}_{1}^{\dagger}\|^{2}_{\textup{F}}\leq{d_{1}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}t^{2},\quad\|\bm{\Omega}_{1}^{\dagger}\|^{2}_{2}\leq{d_{2}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}t^{2},\quad\|\bm{\Omega}_{1}^{\dagger}\|^{2}_{(4)}\leq{d_{3}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}t^{2},$
with the constants
$d_{1}=\frac{3k}{p+1},\quad d_{2}=e^{2}\frac{k+p}{(p+1)^{2}},\quad d_{3}=\sqrt{k}d_{2}.$ (12)
Lemma A.7 implies the following bound for $\mathbb{P}(E_{t}^{c})$:
$\mathbb{P}(E_{t}^{c})\leq 2t^{-(p+1)}+t^{-p}\leq 3t^{-p}.$ (13)
Under the event $E_{t}$, Eq. 11 becomes
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}\mid E_{t}\big{]}^{2}\leq 2\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+2t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}.$ (14)
Conditioning Eq. 10 on $E_{t}$ and combining it with Eq. 14 yields
$\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{2}\leq 4\left(\delta_{k}^{(\textup{F})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\textup{F})}\right)\|\bm{\Sigma}_{2}\|_{\textup{F}}+4t^{2}d_{3}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}+2d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}u^{2}t^{2},$
with probability $\geq 1-e^{-u^{2}/2}$. Similarly to the proof of [20, Thm. 10.8], by the union bound we remove the conditioning by a union bound using Eq. 13 and conclude the proof of Theorem 2.1.
2.2.3 Analysis in the spectral and nuclear norms
The structural bound Eq. 7 for the Nyström method applies to any unitarily invariant matrix norm. Using properties of Gaussian matrices from Appendix A, this allows us to extend the analysis performed in Sections 2.2.1 and 2.2.2 to the spectral and nuclear norms.
Theorem 2.4 (Expectation bound in spectral and nuclear norms).
Consider the setting of Theorem 2.1 with an oversampling parameter $p\geq 2$, it holds that
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{2}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(2)}+3\delta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (15a)
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{*}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}.$ (15b)
Proof 2.5.
We start by proving Eq. 15a and use the structural bound (7) for the Nyström method to obtain
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{2}]\leq\|\bm{\Sigma}_{2}\|_{2}+\mathbb{E}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]=\|\bm{\Sigma}_{2}\|_{2}+\mathbb{E}^{2}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}]^{2}.$
Similarly to the proof of Section 2.2.1, we use the conditional distribution of $\bm{\Omega}_{2}|\bm{\Omega}_{1}$ and the triangle inequality for the $L_{2}$ norm twice, to get
$\mathbb{E}^{2}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\big{]}$ $=\mathbb{E}^{2}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}^{2}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\bm{\Omega}_{1}\bm{\Omega}_{1}^{\dagger}+\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}\big{]}$
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\big{(}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\mid\bm{\Omega}_{1}\big{]}\big{]}\big{)}^{1/2},$
where $\bm{\Psi}$ is an $(n-k)\times(k+p)$ standard Gaussian matrix. Then Eq. 32c leads to
$\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\mid\bm{\Omega}_{1}\big{]}\leq\big{(}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}\|\bm{\Omega}_{1}^{\dagger}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}\big{)}^{2}.$
After taking expectation with respect to $\bm{\Omega}_{1}$ and using the triangle inequality for the $L_{2}$ norm, we have
$\big{(}\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\mathbb{E}_{\bm{\Psi}}\big{[}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\bm{\Psi}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\mid\bm{\Omega}_{1}\big{]}\big{]}\big{)}^{1/2}$ $\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}(\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\big{]})^{1/2}$
$\quad+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}(\mathbb{E}_{\bm{\Omega}_{1}}\big{[}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2}\big{]})^{1/2}.$
We then apply Lemma A.5 to obtain the following inequality
$\mathbb{E}^{2}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}]$ $\leq\frac{\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{\textup{F}}}{\sqrt{p-1}}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}$
$\quad+\frac{e\sqrt{k+p}}{\sqrt{p^{2}-1}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{1/2}\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}$
$\leq\sqrt{\frac{k}{p-1}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\delta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2}}+\sqrt{\frac{e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*}}.$
We conclude the proof of Eq. 15a using the inequality $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$.
The bound for the nuclear norm follows through a similar argument from the structural bound (7) in the nuclear norm:
$\mathbb{E}[\|\bm{A}-\widehat{\bm{A}}\|_{*}]$ $\leq\|\bm{\Sigma}_{2}\|_{*}+\mathbb{E}[\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2}]$
$\leq\|\bm{\Sigma}_{2}\|_{*}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}^{1/2}_{22.1}\|_{\textup{F}}^{2}\frac{\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p-1}$
$\leq\left(1+\frac{k}{p-1}\beta_{k}^{(*)}+\delta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}.$
Theorem 2.6 (Tailbound in spectral and nuclear norms).
Using the notation of Theorem 2.1 with an oversampling parameter $p\geq 4$, and $u,t\geq 1$, it holds that
$\|\bm{A}-\widehat{\bm{A}}\|_{2}$ $\leq\left(1+4\delta_{k}^{(2)}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+4d_{2}t^{2}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$ (16a)
$\|\bm{A}-\widehat{\bm{A}}\|_{*}$ $\leq\left(1+2\delta_{k}^{(*)}+d_{1}t^{2}\beta_{k}^{(*)}\right)\|\bm{\Sigma}_{2}\|_{*}+2d_{2}t^{2}u^{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{2},$ (16b)
where each inequality holds with probability $\geq 1-2t^{-p}-e^{u^{2}/2}$.
Proof 2.7.
We begin by deriving the tailbound Eq. 16a in the spectral norm. Similar to Section 2.2.2, we process the second term of the structural bound Eq. 7 with the concentration inequality of Lemma 2.2 in the spectral norm:
$\mathbb{P}\left\{\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\geq\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{2}u\mid\bm{\Omega}_{1}\right\}\leq e^{-u^{2}/2}.$ (17)
Using Lemma A.3, it holds that
$\mathbb{E}\big{[}\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}\mid\bm{\Omega}_{1}\big{]}$
$\leq\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{\textup{F}}\|\bm{\Omega}_{1}^{\dagger}\|_{2}+\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}^{1/2}\|_{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}.$
We then consider the probability event $E_{t}$ that the Frobenius and spectral norms of $\bm{\Omega}_{1}^{\dagger}$ are well controlled:
$E_{t}=\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}\leq\sqrt{d_{1}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}t,\quad\|\bm{\Omega}_{1}^{\dagger}\|_{2}\leq\sqrt{d_{2}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}t\right\},$
where $\mathbb{P}(E_{t}^{c})\leq 2t^{-p}$ by Lemma A.7. We conclude the proof of Eq. 16a by conditioning Eq. 17 on $E_{t}$ and using the inequality $\|\bm{x}\|_{1}\leq\sqrt{d}\|\bm{x}\|_{2}$ for $\bm{x}\in\mathbb{R}^{d}$ to obtain:
$\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}\leq\left(4\delta_{k}^{(2)}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(2)}\right)\|\bm{\Sigma}_{2}\|_{2}+4d_{2}t^{2}\beta_{k}^{(*)}\|\bm{\Sigma}_{2}\|_{*},$
with probability $\geq 1-2t^{-p}-e^{-u^{2}/2}$.
The proof of the tail bound for the nuclear norm Eq. 16b follows from a similar argument.
Remark 2.8 (Connection with the randomized SVD).
Consider an arbitrary matrix $\bm{B}\in\mathbb{R}^{m\times n}$ with singular value decomposition
$\bm{B}=\begin{bmatrix}\bm{W}_{1}&\bm{W}_{2}\end{bmatrix}\begin{bmatrix}\bm{S}_{1}&\bm{0}\\ \bm{0}&\bm{S}_{2}\end{bmatrix}\begin{bmatrix}\bm{U}_{1}^{*}\\ \bm{U}_{2}^{*}\end{bmatrix},$
where $\bm{W}_{1}\in\mathbb{R}^{m\times k},\bm{U}_{1}\in\mathbb{R}^{n\times k}$, and $\bm{S}_{1}\in\mathbb{R}^{k\times k}$ is a diagonal matrix containing the $k$ largest singular values of $\bm{B}$. Letting $\bm{Q}$ be an orthonormal basis for $\operatorname{range}(\bm{B}\bm{\Omega})$, then $\bm{Q}\bm{Q}^{*}\bm{B}$ is the approximation attained by the (basic) randomized SVD. Setting $\bm{A}=\bm{B}^{*}\bm{B}$ and $\widehat{\bm{A}}=\bm{A}\bm{\Omega}(\bm{\Omega}^{*}\bm{A}\bm{\Omega})^{\dagger}\bm{\Omega}^{*}\bm{A}$ we have for any $s\geq 1$ that
$\|\bm{B}-\bm{Q}\bm{Q}^{*}\bm{B}\|_{(2s)}^{2}=\|\bm{A}-\widehat{\bm{A}}\|_{(s)}.$
Hence, obtaining a bound for the randomized SVD applied to $\bm{B}$ is equivalent to obtaining a bound for the Nyström approximation on $\bm{A}$; similar observations have been made in [19, 38]. Therefore, one can apply the results obtained earlier in this section to derive error bounds for the randomized SVD. As an example, Theorem 2.4 implies that
$\mathbb{E}\|\bm{B}-\bm{Q}\bm{Q}^{*}\bm{B}\|_{\textup{F}}^{2}\leq\left(1+\frac{k}{p-1}\tilde{\beta}_{k}+\tilde{\delta}_{k}\right)\|\bm{S}_{2}\|_{\textup{F}}^{2},$
where $\tilde{\beta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{22.1}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}$ and $\tilde{\delta}_{k}=\operatorname{tr}(\bm{S}_{2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{S}_{2})\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}/\|\bm{S}_{2}\|_{\textup{F}}^{2}$. This bound coincides with the standard randomized SVD bound [20, Thm. 10.5] when $\bm{K}=\bm{I}$, unlike the bound proved in [7, Thm. 2].
3 Nyström approximation in infinite dimensions
This section presents an infinite-dimensional extension of the randomized Nyström approximation. We begin by briefly introducing the concepts of quasimatrices, Hilbert–Schmidt operators, and Gaussian processes, which will be useful to generalize the bounds of Section 2 to operators between function spaces. While we consider operators between square-integrable functions as the most important application, the results of this section extend naturally to other separable Hilbert spaces following the theory in [22].
3.1 Quasimatrices
For a bounded domain $D\subset\mathbb{R}^{d}$, $d\geq 1$, we consider the Hilbert space $L^{2}(D)$ of square-integrable functions. Quasimatrices are a convenient way to represent and work with collections of functions or more, generally, elements of infinite-dimensional vector spaces; see, e.g., [37]. In particular, a function $Y:\mathbb{R}^{m}\to L^{2}(D)$ is expressed as the quasimatrix
$Y=\begin{bmatrix}y_{1}&\cdots&y_{m}\end{bmatrix},\quad y_{i}\in L^{2}(D).$
[…]
$Y^{*}Z=\begin{bmatrix}\langle y_{1},z_{1}\rangle&\cdots&\langle y_{1},z_{\ell}\rangle\\ \vdots&\ddots&\vdots\\ \langle y_{m},z_{1}\rangle&\cdots&\langle y_{m},z_{\ell}\rangle\end{bmatrix}.$
3.2 Non-negative self-adjoint trace-class operators
We consider a non-negative self-adjoint trace-class operator $\mathcal{A}:L^{2}(D)\to L^{2}(D)$, i.e., it holds that $\langle\mathcal{A}f,f\rangle\geq 0$ and $\langle\mathcal{A}f,g\rangle=\langle f,\mathcal{A}g\rangle$ for every $f,g\in L^{2}(D)$, and the trace norm [22, Def. 4.5.1] is finite:
$\|\mathcal{A}\|_{\mathrm{Tr}}:=\sum_{j=1}^{\infty}\langle\mathcal{A}e_{j},e_{j}\rangle<\infty,$
for any orthonormal basis $\{e_{j}\}_{j}$ of $L^{2}(D)$. Non-negative self-adjoint trace-class operators are Hilbert–Schmidt operators [22, Thm. 4.5.2], and therefore admit an eigenvalue decomposition of the form [22, Thm. 4.3.1]:
$\mathcal{A}=\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}>0\end{subarray}}^{\infty}\sigma_{j}\langle u_{j},\cdot\rangle u_{j},$ (18)
where $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$ are the eigenvalues of $\mathcal{A}$, and $\{u_{j}\}_{j}$ are orthonormal eigenfunctions. The eigenvalues allow us to express the trace, Hilbert–Schmidt, and operator norms of $\mathcal{A}$ as
$\|\mathcal{A}\|_{\mathrm{Tr}}=\sum_{j=1}^{\infty}\sigma_{j},\quad\|\mathcal{A}\|_{\mathrm{HS}}=\Big{(}\sum_{j=1}^{\infty}\sigma_{j}^{2}\Big{)}^{1/2},\quad\|\mathcal{A}\|_{\mathrm{op}}=\sigma_{1},$
which are infinite-dimensional analogs of the nuclear, Frobenius, and spectral norms discussed in Section 2. Furthermore, we introduce the operator $U:\ell^{2}\to L^{2}(D)$ defined by $Uf=\sum_{i=1}^{\infty}f_{i}u_{i}$ for any $f$ in $\ell^{2}$, the space of square-summable sequences (indexed by positive integers). Then, for a given rank $k\geq 1$, the quasimatrix $U_{1}:\mathbb{R}^{k}\to L^{2}(D)$ contains the first $k$ eigenfunctions of $\mathcal{A}$, and the quasimatrix $U_{2}:\ell^{2}\to L^{2}(D)$ contains the remaining eigenfunctions. Finally, we introduce the diagonal matrix and quasimatrices $\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k})$ and $\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\sigma_{k+2},\ldots)$, respectively, which contain the eigenvalues of $\mathcal{A}$ in descending order.
3.3 Gaussian processes
Let $K:D\times D\mapsto\mathbb{R}$ be a continuous symmetric positive semi-definite kernel, with the following eigenvalue decomposition [22, Thm. 4.6.5]:
[…]
where the sum converges absolutely and uniformly on $D$ [29]. Here, $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq 0$ are the eigenvalues of the integral operator $\mathcal{K}$ induced by $K$:
$\mathcal{K}[f](x)=\int_{D}K(x,y)f(y)\,\mathrm{d}y,\quad f\in L^{2}(D),\,x\in D,$
and $\{\psi_{j}\}_{j}$ are the corresponding orthonormal eigenfunctions of $\mathcal{K}$ in $L^{2}(D)$. In the following, we assume that $\mathcal{K}$ is trace-class, i.e., $\sum_{i=1}^{\infty}\lambda_{i}<\infty$. A Gaussian process (GP) $\omega\sim\mathcal{GP}(0,K)$ is a stochastic process such that for any finite collection of points $x_{1},\ldots,x_{m}\in D$, the vector $\begin{bmatrix}\omega(x_{1})&\cdots&\omega(x_{m})\end{bmatrix}^{\top}$ follows a multivariate normal distribution with mean $\bm{0}\in\mathbb{R}^{m}$ and covariance matrix $\bm{K}=(K(x_{i},x_{j}))_{1\leq i,j\leq m}$. The Karhunen–Loève expansion [1, Ch.3.2] of $\omega$ is given by
$\omega\sim\sum_{k=1}^{\infty}\lambda_{k}^{1/2}\zeta_{k}\psi_{k},$ (19)
where $\zeta_{1},\zeta_{2},\ldots\sim\mathcal{N}(0,1)$ are mutually independent. With probability one, a realization of $\omega$ is in $L^{2}(D)$ [22, Thm. 7.2.5]. Recalling that $\{u_{j}\}_{j}$ denote the eigenfunctions of $\mathcal{A}$, see (18), we define the function $\widetilde{\bm{K}}:\mathbb{N}^{*}\times\mathbb{N}^{*}\to\mathbb{R}$ elementwise as
$\widetilde{\bm{K}}(i,j)=\langle u_{i},\mathcal{K}u_{j}\rangle=\sum_{k=1}^{\infty}\lambda_{k}\langle\psi_{k},u_{i}\rangle\langle\psi_{k},u_{j}\rangle,\quad i,j\in\mathbb{N}^{*},$
which is bounded by $\|\mathcal{K}\|_{\mathrm{Tr}}$ using the Cauchy–Schwarz inequality. We also define the following restrictions of $\widetilde{\bm{K}}$ to different sets of indices:
$\widetilde{\bm{K}}_{11}=\widetilde{\bm{K}}_{\llbracket 1,k\rrbracket\times\llbracket 1,k\rrbracket},\quad\widetilde{\bm{K}}_{21}=\widetilde{\bm{K}}_{\llbracket k+1,\infty)\times\llbracket 1,k\rrbracket},\quad\widetilde{\bm{K}}_{22}=\widetilde{\bm{K}}_{\llbracket k+1,\infty)\times\llbracket k+1,\infty)}.$
Then, $\widetilde{\bm{K}}_{11}$ defines a $k\times k$ matrix, which we assume to be of rank $k$ in the rest of this section. In terms of the quasimatrices $U_{1},U_{2}$ containing the eigenfunctions defined above, we can write
$\widetilde{\bm{K}}_{11}=U_{1}^{*}\mathcal{K}U_{1},\quad\widetilde{\bm{K}}_{21}=U_{2}^{*}\mathcal{K}U_{1},\quad\widetilde{\bm{K}}_{22}=U_{2}^{*}\mathcal{K}U_{2},$
in analogy to Eq. 3.
3.4 Infinite-dimensional extension of the Nyström approximation
We are now ready to present the infinite-dimensional extension of the Nyström approximation. Let $k$ be a target rank, $p$ be an oversampling parameter, and $\Omega=\begin{bmatrix}\omega_{1}&\cdots&\omega_{k+p}\end{bmatrix}$ be a random quasimatrix with $k+p$ columns, whose columns are i.i.d. from $\mathcal{GP}(0,K)$. The Nyström approximation $\widehat{\mathcal{A}}$ to $\mathcal{A}$ is defined as
$\widehat{\mathcal{A}}:=\mathcal{A}\Omega(\Omega^{*}\mathcal{A}\Omega)^{\dagger}(\mathcal{A}\Omega)^{*}.$ (20)
Assuming that the realization of $\omega_{i}$ is in $L^{2}(D)$, which holds with probability $1$, the Nyström approximation is an operator $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ of rank at most $k+p$ with the explicit representation
$\widehat{\mathcal{A}}[f]=\sum_{i,j=1}^{k+p}\mathcal{A}\omega_{i}\big{[}(\Omega^{*}\mathcal{A}\Omega)^{\dagger}\big{]}_{ij}\langle\omega_{j},\mathcal{A}f\rangle.$
As in the finite-dimensional case, the error bounds for the infinite-dimensional analog of the Nyström approximation depend on the prior information of eigenvectors contained in $\mathcal{K}$, which is measured by the following two quantities:
$\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},$ (21)
where $\xi\in\{\mathrm{HS},\mathrm{Tr},\mathrm{op}\}$. These quantities are the infinite-dimensional analogs of Eq. 4.
For each $1\leq j\leq k+p$, we consider the stochastic process $\bm{\omega}_{j}=\{\langle u_{i},\omega_{j}\rangle,\,i\in\mathbb{N}^{*}\}$ whose trajectories are in $\ell^{2}$, and denote by $\bm{\Omega}=U^{*}\Omega=\begin{bmatrix}\bm{\omega}_{1}&\cdots&\bm{\omega}_{k+p}\end{bmatrix}$ the random quasimatrix whose columns are i.i.d. from $\mathcal{GP}(0,\widetilde{\bm{K}})$. Then, we introduce the random $k\times(k+p)$ matrix $\bm{\Omega}_{1}$ as
$\bm{\Omega}_{1}\coloneqq U_{1}^{*}\Omega=\begin{bmatrix}\langle u_{1},\omega_{1}\rangle&\cdots&\langle u_{1},\omega_{k+p}\rangle\\ \vdots&\ddots&\vdots\\ \langle u_{k},\omega_{1}\rangle&\cdots&\langle u_{k},\omega_{k+p}\rangle\end{bmatrix},\quad\omega_{j}\sim\mathcal{GP}(0,K),\quad 1\leq j\leq k+p,$
whose columns are i.i.d. from $\mathcal{N}(0,\widetilde{\bm{K}}_{11})$ [8, Lem. 1]. Since we assume that $\operatorname{rank}(\widetilde{\bm{K}}_{11})=k$ we know that $\bm{\Omega}_{1}$ has full row-rank with probability one. Therefore, $\bm{\Omega}_{1}$ has almost surely a right inverse $\bm{\Omega}_{1}^{\dagger}=\bm{\Omega}_{1}^{*}(\bm{\Omega}_{1}\bm{\Omega}_{1}^{*})^{-1}$. We define $\bm{\Omega}_{2}=U_{2}^{*}\Omega$ similarly. Note that the $i^{\text{th}}$ entry of the columns of $\bm{\Omega}_{2}$ are distributed as $\mathcal{N}(0,(\widetilde{\bm{K}}_{22})_{ii})$ and the covariance between the $i^{\text{th}}$ and $j^{\text{th}}$ entry is $(\widetilde{\bm{K}}_{22})_{ij}$ [8, Sec. 3.3].
3.4.1 Structural bound
In this section, we prove an infinite-dimensional analog of the structural bound (7). We begin by stating some basic but useful results on norms of finite sections of $\ell^{2}$ operators.
Lemma 3.1.
Let $\bm{B},\bm{C}:\ell^{2}\to\ell^{2}$ be non-negative self-adjoint trace-class operators such that $\bm{B}-\bm{C}$ is non-negative. For $n\in\mathbb{N}^{*}$, consider the restriction $\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}$ of $\bm{B}$ to its first $n$ rows and columns. Then, for $\xi\in\{\mathrm{Tr},\mathrm{HS},\mathrm{op}\}$, we have
$\lim_{n\rightarrow\infty}\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}$ $=\|\bm{B}\|_{\xi},$ (22a)
$\|\bm{C}\|_{\xi}\leq\|\bm{B}\|_{\xi}$ $\leq\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}+\|\bm{B}_{\llbracket n+1,\infty)\times\llbracket n+1,\infty)}\|_{\xi}.$ (22b)
For $\xi\in\{\mathrm{HS},\mathrm{Tr}\}$, the property (22a) follows from the absolute convergence of the involved series. For $\xi=\mathrm{op}$, the triangle inequality yields
$|\|\bm{B}\|_{\mathrm{op}}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{op}}|\leq\sqrt{\|\bm{B}\|_{\mathrm{HS}}^{2}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{HS}}^{2}}\rightarrow 0,\quad\text{as }n\to\infty.$
Eq. Eq. 22b is an infinite-dimensional analog of [25, Thm. 2.1] and [32, Lem. 3.1], which can be proven using Eq. 22a. We can now proceed with an infinite-dimensional analog of the finite-dimensional structural bound Eq. 7.
Lemma 3.2 (Infinite-dimensional structural bound).
Let $\mathcal{A}:L^{2}(D)\to L^{2}(D)$ be a self-adjoint non-negative trace-class operator, $k,p\geq 1$, and $\Omega:\mathbb{R}^{k+p}\to L^{2}(D)$ be a quasimatrix with $k+p$ columns such that the matrix $\bm{\Omega}_{1}=U_{1}^{*}\Omega\in\mathbb{R}^{k\times(k+p)}$ is full rank. Then,
$\|\mathcal{A}-\mathcal{A}\Omega(\Omega^{*}\mathcal{A}\Omega)^{\dagger}(\mathcal{A}\Omega)^{*}\|_{\xi}\leq\|\bm{\Sigma}_{2}\|_{\xi}+\|(\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger})^{*}\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{\xi},$
where $\xi\in\{\mathrm{op},\mathrm{HS},\mathrm{Tr}\}$.
Proof 3.3.
Let $\bm{\Omega}=U^{*}\Omega$ be defined as in Section 3.4 and $\mathcal{P}_{U}=UU^{*}:L^{2}(D)\to L^{2}(D)$ denote the orthogonal projection onto the range of $\mathcal{A}$. Since $\mathcal{A}$ is a self-adjoint operator, we have $\mathcal{A}=\mathcal{P}_{U}\mathcal{A}=\mathcal{A}\mathcal{P}_{U}=\mathcal{P}_{U}\mathcal{A}\mathcal{P}_{U}$. Therefore,
$\|\mathcal{A}-\mathcal{A}\Omega(\Omega^{*}\mathcal{A}\Omega)^{\dagger}(\mathcal{A}\Omega)^{*}\|_{\xi}=\|\mathcal{P}_{U}\mathcal{A}\mathcal{P}_{U}-\mathcal{P}_{U}\mathcal{A}\mathcal{P}_{U}\Omega(\Omega^{*}\mathcal{P}_{U}\mathcal{A}\mathcal{P}_{U}\Omega)^{\dagger}(\mathcal{P}_{U}\mathcal{A}\mathcal{P}_{U}\Omega)^{*}\|_{\xi}$
$\quad\qquad=\|U(U^{*}\mathcal{A}U-U^{*}\mathcal{A}U\bm{\Omega}(\bm{\Omega}^{*}U^{*}\mathcal{A}U\bm{\Omega})^{\dagger}(U^{*}\mathcal{A}U\bm{\Omega})^{*})U^{*}\|_{\xi}$
$\quad\qquad=\|U^{*}\mathcal{A}U-U^{*}\mathcal{A}U\bm{\Omega}(\bm{\Omega}^{*}U^{*}\mathcal{A}U\bm{\Omega})^{\dagger}(U^{*}\mathcal{A}U\bm{\Omega})^{*}\|_{\xi}=\|\bm{\Sigma}-\bm{\Sigma}\bm{\Omega}(\bm{\Omega}^{*}\bm{\Sigma}\bm{\Omega})^{\dagger}(\bm{\Sigma}\bm{\Omega})^{*}\|_{\xi},$
where the third equality follows from the unitary invariance of the norms and the fourth equality is due to the relation $U^{*}\mathcal{A}U=\bm{\Sigma}$. As in the finite-dimensional case, the rest of the proof follows the argument of the proof of [32, Lem. 3.13] for the operator monotone function $f:x\mapsto x$ and $q=1$, now using the fact that the inequalities in Eq. 22b are infinite-dimensional analogs of [32, Lems. A.1 and 3.1] with $f:x\mapsto x$ and $\|\cdot\|\coloneqq\|\cdot\|_{\xi}$.
3.5 Probabilistic bounds
With the structural bound in place, we proceed to derive probabilistic bounds for the infinite-dimensional Nyström approximation (20).
Theorem 3.4 (Infinite-dimensional Nyström approximation).
Let $\mathcal{A}:L^{2}(D)\to L^{2}(D)$ be a non-negative self-adjoint trace-class operator, $2\leq k\leq\operatorname{rank}(\mathcal{A})$ be a target rank, and $p\geq 4$ be an oversampling parameter. Let $\Omega$ be a quasimatrix with $k+p$ columns i.i.d. from $\mathcal{GP}(0,K)$, with a kernel $K$ such that the matrix $\widetilde{\bm{K}}_{11}$ defined in Eq. 3 is invertible. Then, the Nyström approximation $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ defined in Eq. 20 satisfies
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}]$ $\leq\left(1+2\delta_{k}^{(\mathrm{HS})}+2\sqrt{c_{1}}\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+2\sqrt{c_{2}}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23a)
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(\mathrm{op})}+3\delta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23b)
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(\mathrm{Tr})}+\delta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23c)
where $c_{1}=\mathcal{O}(k^{2}/p^{2})$, $c_{2}=\mathcal{O}(k^{2}/p^{2})$ are the constants defined in Eq. 9. Let $u,t\geq 1$, then
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}$ $\leq\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4\left(\delta_{k}^{(\mathrm{HS})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4t^{2}d_{3}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}$ (24a)
$\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},$
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}$ $\leq\left(1+4\delta_{k}^{(\mathrm{op})}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+4d_{2}t^{2}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (24b)
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}$ $\leq\left(1+2\delta_{k}^{(\mathrm{Tr})}+d_{1}t^{2}\beta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+2d_{2}t^{2}u^{2}\beta_{k}^{(\mathrm{op})}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},$ (24c)
with probability $\geq 1-3t^{-p}-e^{-u^{2}/2}$. Here, $d_{1}=\mathcal{O}(k/p)$, $d_{2}=\mathcal{O}(k/p)$, and $d_{3}=\mathcal{O}(k^{3/2}/p)$ are the constants defined in Eq. 12.
The proof of Theorem 3.4 occupies the rest of this section and follows from a continuity argument on the generalization of the Nyström approximation to correlated Gaussian vectors analyzed earlier in Section 2. Let $n\geq 1$ and $s\in[1,\infty]$, we first define the following random variables:
$X_{s,n}=\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Omega}_{2})_{\llbracket 1,n\rrbracket\times\llbracket 1,k+p\rrbracket}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},\quad X_{s}=\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},$ (25)
where $\|\cdot\|_{(2s)}$ denotes the Schatten-$2s$ norm. We aim to show the convergence of $X_{s,n}$ to $X_{s}$ as $n\to\infty$, and begin with a preliminary result on the finiteness of the expectation of $X_{s}$.
Lemma 3.5 (Expectation of $X_{s}$).
For $s\in[1,\infty]$, let $X_{s}$ be the random variable defined in Eq. 25. Then, if $p\geq 4$, we have $\mathbb{E}_{\Omega}^{2}[X_{s}]<\infty$.
Proof 3.6.
We first notice that $2s\geq 2$ implies
$\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(2s)}^{2}\leq\|\bm{\Sigma}_{2}^{1/2}\|_{\mathrm{op}}^{2}\|\bm{\Omega}_{2}\|_{\mathrm{HS}}^{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2}.$
Noting that $\bm{\Omega}_{1}$, $\bm{\Omega}_{2}$ are not independent, we need to establish $\mathbb{E}[\|\bm{\Omega}_{2}\|_{\mathrm{HS}}^{4}]<\infty$ and $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]<\infty$ in order to conclude the result from Hölder’s inequality. First, Lemma A.5 ensures that $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]<\infty$ since $\bm{\Omega}_{1}$ is a $k\times(k+p)$ matrix whose columns are i.i.d. from $\mathcal{N}(\bm{0},\widetilde{\bm{K}}_{11})$ [8, Lem. 1]. By the Karhunen–Loève expansion (19), an arbitrary column $\omega$ of $\Omega$ satisfies
$\mathbb{E}\big{[}\|\omega\|^{4}_{L^{2}(D)}\big{]}$ =$\sum_{i,j=1}^{\infty}\mathbb{E}[\zeta^{2}_{i}\zeta^{2}_{j}]\lambda_{i}\lambda_{j}=\sum_{i=1}^{\infty}\mathbb{E}[\zeta^{4}_{i}]\lambda_{i}^{2}+\sum_{i\not=j}\mathbb{E}[\zeta_{i}^{2}\zeta_{j}^{2}]\lambda_{i}\lambda_{j}$
=$3\sum_{i=1}^{\infty}\lambda_{i}^{2}+\sum_{i\not=j}\lambda_{i}\lambda_{j}\leq 3\|\mathcal{K}\|_{\mathrm{Tr}}^{2}<\infty.$
In turn,
$\mathbb{E}\|\bm{\Omega}_{2}\|_{\mathrm{HS}}^{4}\leq\mathbb{E}\|\Omega\|_{\mathrm{HS}}^{4}\leq\mathbb{E}\big{(}\|\omega_{1}\|^{2}_{L^{2}(D)}+\cdots+\|\omega_{k+p}\|^{2}_{L^{2}(D)}\big{)}^{2}\leq(k+p)^{2}\mathbb{E}\big{[}\|\omega\|^{4}_{L^{2}(D)}\big{]}<\infty.$
We proceed by showing that $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s}-X_{s,n}]=0$.
Lemma 3.7 (Convergence of $X_{s,n}$ to $X_{s}$).
For $s\in[1,\infty]$, let $X_{s}$, $X_{s,n}$ be the random variables defined in Eq. 25 for $n\geq 1$. Then, if $p\geq 4$, we have $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s}-X_{s,n}]=0$.
Proof 3.8.
For $n\geq 1$, we define the quasimatrix $\bm{\Omega}_{2}^{(n)}$ whose first $n$ rows are equal to the first $n$ rows of $\bm{\Omega}_{2}$ and the remaining rows are zero. Then,
$X_{s,n}=\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Omega}_{2})_{\llbracket 1,n\rrbracket\times\llbracket 1,k+p\rrbracket}\bm{\Omega}_{1}^{\dagger}\|_{(2s)}=\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}^{(n)}\bm{\Omega}_{1}^{\dagger}\|_{(2s)}.$
Combining the triangle inequality and sub-multiplicativity of the Schatten-$s$ norm, and using the fact that $2s\geq 2$, we have
$(X_{s}-X_{s,n})^{2}\leq\|\bm{\Sigma}_{2}^{1/2}\|_{\mathrm{op}}^{2}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{2}\|\bm{\Omega}_{2}-\bm{\Omega}_{2}^{(n)}\|_{\mathrm{HS}}^{2}.$
[…]
$\mathbb{E}^{4}[\|\bm{\Omega}_{2}-\bm{\Omega}_{2}^{(n)}\|_{\mathrm{HS}}]$ $\leq\mathbb{E}^{4}\left[\sum_{i=1}^{k+p}\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(n)}\|_{2}\right]\leq\sum_{i=1}^{k+p}\mathbb{E}^{4}\left[\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(n)}\|_{2}\right]$
$=(k+p)\mathbb{E}^{4}\left[\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}\right],$
since the columns of $\bm{\Omega}_{2}$ are identically distributed.
We are now going to verify that
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{4}\right]=0.$
Following [8, Sec. 3.3], we know that the entries of $\bm{\omega}_{1}$ satisfy $(\bm{\omega}_{1})_{i}\sim\mathcal{N}(0,(\widetilde{\bm{K}}_{22})_{ii})$ for $i\in\mathbb{N}^{*}$. Let $Y=\sum_{i=n+1}^{\infty}(\bm{\omega}_{1})_{i}^{2}=\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{2}$. Combining the non-negativity of the summands and the Fubini–Tonelli theorem, we can interchange the summation and expectation to obtain
$\mathbb{E}\left[\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{4}\right]$ $=\mathbb{E}[Y^{2}]=\mathbb{E}\left[\sum_{i=n+1}^{\infty}(\bm{\omega}_{1})_{i}^{2}Y\right]=\sum_{i=n+1}^{\infty}\mathbb{E}\left[(\bm{\omega}_{1})_{i}^{2}Y\right]$
$\leq\sum_{i=n+1}^{\infty}\sqrt{\mathbb{E}[(\bm{\omega}_{1})_{i}^{4}]}\sqrt{\mathbb{E}[Y^{2}]},$
where the last inequality follows from the Cauchy–Schwarz inequality. Hence,
$\mathbb{E}\left[\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{4}\right]\leq\left(\sum_{i=n+1}^{\infty}\sqrt{\mathbb{E}[(\bm{\omega}_{1})_{i}^{4}]}\right)^{2}=3\left(\sum_{i=n+1}^{\infty}(\widetilde{\bm{K}}_{22})_{ii}\right)^{2}\rightarrow 0,\quad\text{as }n\rightarrow\infty,$
since $\operatorname{tr}(\widetilde{\bm{K}}_{22})\leq\operatorname{tr}(\mathcal{K})<\infty$.
Combining Lemmas 3.5 and 3.7, we obtain that $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s,n}]=\mathbb{E}^{2}[X_{s}]$. Hence, if we have a family of bounds $\mathbb{E}^{2}[X_{s,n}]\leq y_{s,n}$ with $y_{s,n}\to y_{s}$ as $n\to\infty$, then $\mathbb{E}^{2}[X_{s}]\leq y_{s}$. Furthermore, combining the continuous mapping theorem and the fact that $L^{2}$ convergence implies convergence in distribution, for any positive sequence $z_{s,n}\to z_{s}<\infty$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}^{2}>z_{s,n})=\mathbb{P}(X_{s}^{2}>z_{s})$. We can now proceed with the proof of Theorem 3.4, which uses results from Section 2 to derive expressions for $y_{s,n}$ and $z_{s,n}$ and show that, for $s\in\{2,\infty,1\}$, they converge to the right-hand sides of Theorems 3.4 and 3.4.
Proof 3.9 (Proof of Theorem 3.4).
Let $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+y_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+y_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+y_{1}$ be the respective right-hand sides in Theorem 3.4, and $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+z_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+z_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+z_{1}$ be the right-hand sides in Theorem 3.4. For $s\in\{2,\infty,1\}$, we aim to prove that $\mathbb{E}[X_{s}^{2}]\leq y_{s}$ and $\mathbb{P}(X_{s}^{2}\geq z_{s})\leq 3t^{-p}+e^{-u^{2}/2}$. Following Theorem 2.1, we have
$\mathbb{E}[X_{2,n}^{2}]$ $\leq 2\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+2\sqrt{c_{1}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+2\sqrt{c_{2}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}:=y_{2,n},$
and with probability greater than $1-3^{-p}-e^{-u^{2}/2}$, we have
$X_{2,n}^{2}$ $\geq 4\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+4t^{2}(d_{1}+d_{3})\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+4t^{2}d_{3}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}$
$\quad+2t^{2}u^{2}d_{2}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{2}:=z_{2,n}.$
Following Theorems 2.4 and 2.6 we know that
$\mathbb{E}[X_{\infty,n}]\leq y_{\infty,n},$ $\mathbb{E}[X_{1,n}]\leq y_{1,n},$
$\mathbb{P}\left(X_{\infty,n}>z_{\infty,n}\right)\leq 3^{-p}+e^{-u^{2}/2},$ $\mathbb{P}\left(X_{1,n}>z_{1,n}\right)\leq 3^{-p}+e^{-u^{2}/2},$
where $y_{\infty,n}$, $y_{1,n}$, $z_{\infty,n}$, and $z_{1,n}$ can be defined analogously to $y_{2,n}$ and $z_{2,n}$ using Theorems 2.4 and 2.6. Moreover, Lemma 3.7 implies that $\lim_{n\rightarrow\infty}\mathbb{E}[X_{s,n}^{2}]=\mathbb{E}[X_{s}^{2}]$ and convergence in distribution, which implies $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}>z_{s,n})=\mathbb{P}(X_{s}>\lim_{n\rightarrow\infty}z_{s,n})$. Hence, it is sufficient to show that $\lim_{n\rightarrow\infty}y_{s,n}=y_{s}$ and $\lim_{n\rightarrow\infty}z_{s,n}=z_{s}$ for $s\in\{2,\infty,1\}$. For this purpose, it is sufficient to show that
$\lim_{n\rightarrow\infty}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}=\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi},$
$\lim_{n\rightarrow\infty}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}=$ (26)
$\hskip 273.18271pt\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi},$
for $\xi\in\{\mathrm{HS},\mathrm{Tr},\mathrm{op}\}$, as inserting the definitions of $\beta_{k}^{(\xi)}$ and $\delta_{k}^{(\xi)}$ would then give the result. Note that since $\bm{\Sigma}_{2}$ is diagonal we have
$(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}=(\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket},$
$(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}=$
$\hskip 260.17464pt(\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}.$
Since $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}$ and $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}$ are non-negative trace-class operators, applying (22a) yields (26), as desired.
4 Numerical experiments
In this section, we test the infinite-dimensional Nyström approximation proposed in this work to compute low-rank approximations to different integral kernel operators. Algorithm 1 presents the pseudocode of a variant of [28, Alg. 16] for non-negative self-adjoint trace-class operators. We emphasize that the orthonormalization steps in line 2 and the shift in line 5 improve the numerical stability of the algorithm; see e.g. [39]. Moreover, our implementation of the Nyström approximation computes the Cholesky factorization of the symmetric part of $\bm{\Omega}^{*}\bm{Y}_{\nu}$ in line 6. Finally, we note that in exact arithmetic, with $\nu=0$ in line 4, the approximation returned by Algorithm 1 is mathematically equivalent to Eq. 20.
The aim of our numerical experiments is to validate Algorithm 1. For this purpose, we consider an interval $D=[a,b]$ and carry out all operations on functions on $D$ (approximately) using the Chebfun software package [15], which is an open-source MATLAB package containing algorithms for performing numerical linear algebra operations on functions, such as the QR decomposition [36, 37]. In all experiments, we select the squared-exponential kernel:
$K(x,y)=\exp\left(-\frac{2(x-y)^{2}}{(b-a)^{2}\ell^{2}}\right),\quad x,y\in[a,b],$
as covariance kernel associated with the random field $\Omega$ in Algorithm 1. We vary the length-scale parameter $\ell$ across three values: $\ell=1,0.1,0.01$. A smaller value for $\ell$ results in smoother Gaussian processes that are more biased towards certain spatial directions. Conversely, a larger value of $\ell$ results in rougher Gaussian processes that are less spatially biased. We vary $\ell$ to investigate the effect of the smoothness of the Gaussian process on the result of the low-rank approximation.
Algorithm 1 Infinite-dimensional Nyström approximation
input: Non-negative self-adjoint trace-class $\mathcal{A}:L^{2}(D)\to L^{2}(D)$, covariance kernel $K:D\times D\to\mathbb{R}$. Target rank $k$, oversampling parameter $p$.
output: Rank $k+p$ Nyström approximation $\widehat{\mathcal{A}}$ to $\mathcal{A}$ in factored form.
1:Draw a random quasimatrix $\Omega=\begin{bmatrix}\omega_{1}&\cdots&\omega_{k+p}\end{bmatrix}$ with columns i.i.d. from $\mathcal{GP}(0,K)$.
2:Orthonormalize columns of $\Omega$: $Q=\text{orth}(\Omega)=\begin{bmatrix}q_{1}&\cdots&q_{k+p}\end{bmatrix}$.
3:Apply operator $\mathcal{A}$ to $Q$: $Y=\mathcal{A}Q=\begin{bmatrix}\mathcal{A}q_{1}&\cdots&\mathcal{A}q_{k+p}\end{bmatrix}$.
4:$\nu=\epsilon\|Y\|_{\mathrm{HS}}$ where $\epsilon$ is equal to the machine precision.
5:Compute shifted $Y_{\nu}=Y+\nu\Omega$.
6:Compute Cholesky factorization of $\bm{\Omega}^{*}\bm{Y}_{\nu}=\bm{R}^{T}\bm{R}$. $\triangleright$ First compute the symmetric part of $\bm{\Omega}^{*}\bm{Y}_{\nu}$ if needed.
7:Perform a triangular solve to compute $B=Y_{\nu}\bm{R}^{-1}$.
8:Compute the Hilbert-Schmidt decomposition of $B=\widehat{U}\bm{S}\bm{V}^{*}$.
9:Remove shift $\widehat{\bm{\Sigma}}=\max\{\bm{S}^{2}-\nu\bm{I},0\}$, where the maximum is taken entry-wise.
10:return $\widehat{\mathcal{A}}=\widehat{U}\widehat{\bm{\Sigma}}\widehat{U}^{*}$ in factored form.
4.1 A pretty function
In this first example, we compute an approximation to the integral operator defined by the kernel [35]
$G(x,y)=\frac{1}{1+100(x^{2}-y^{2})^{2}},\quad x,y\in[-1,1],$ (27)
using Algorithm 1. We vary the number of sketches $k$ (i.e., target rank) and report in Fig. 1 the relative error in the Hilbert-Schmidt norm or, equivalently the $L^{2}$ norm of $G-G_{k}$, where $G_{k}$ is the low-rank Nyström approximant. As can be seen from the figures, setting $\ell=1$ yields a poor approximation to the kernel where the error stagnates for $k>15$, and $\ell=0.01$ is better to obtain high-accuracy approximations.
[…]
Figure 1: (a) Exact kernel defined by Eq. 27 along with convergence of the Nyström approximation for different values of $\ell$ (b). (c)-(e) Rank-$40$ Nyström approximations of the kernel for $\ell=1,0.1,0.01$, respectively.
4.2 Matérn Kernels
In this second example, we approximate the integral operator defined by the Matérn-$1/2,3/2,$ and $5/2$ kernels [34, Chapt. 4]
$G_{1/2}(x,y)=\exp(-|x-y|),\quad x,y\in[-1,1];$ (28)
$G_{3/2}(x,y)=(1+\sqrt{3}|x-y|)\exp\left(-\sqrt{3}|x-y|\right),\quad x,y\in[-1,1];$ (29)
$G_{5/2}(x,y)=\left(1+\sqrt{5}|x-y|+\frac{5}{3}(x-y)^{2}\right)\exp\left(-\sqrt{5}|x-y|\right),\quad x,y\in[-1,1].$ (30)
The Matérn class is an important class of covariance kernels, frequently appearing in machine learning. The parameter $\nu=1/2,3/2,5/2$ determines the spectral decay of the kernel and thus the smoothness of the Gaussian process, with higher $\nu$ implying faster decay and smoother Gaussian processes. The results are presented in Fig. 2 for $G_{1/2},G_{3/2},$ and $G_{5/2}$, respectively.
[…]
Figure 2: The contour plots show the exact kernel defined by (28) and the approximated kernels using rank $100$ Nyström approximations. The error plots show the relative error in the Hilbert-Schmidt norm. Optimal denotes the best low-rank approximation error.
4.3 Green’s function for an elliptic differential operator
In this example we consider the operator $\mathcal{A}_{\lambda}$ so that $u=\mathcal{A}_{\lambda}f$ solves the equation
$-\Delta u(\bm{x})+\lambda u(\bm{x})=f(\bm{x}),\quad\bm{x}\in[0,2\pi]^{d},$
where $\lambda\geq 0$, with Dirichlet boundary conditions and $d\in\{1,2,3\}$ the spatial dimension. Let $\mathcal{A}_{0}$ have Hilbert-Schmidt decomposition
$[\mathcal{A}_{0}f](\bm{x})=\sum\limits_{i=1}^{\infty}\sigma_{i}u_{i}(\bm{x})\langle u_{i},f\rangle,$
[…]
$[\mathcal{A}_{\lambda}f](\bm{x})=\sum\limits_{i=1}^{\infty}g_{\lambda}(\sigma_{i})u_{i}(\bm{x})\langle u_{i},f\rangle=[g_{\lambda}(\mathcal{A}_{0})f](\bm{x}),\quad g_{\lambda}(\sigma)=\frac{\sigma}{\lambda\sigma+1}.$
For each $\lambda\geq 0$ the function $g_{\lambda}$ is operator monotone [5, Sec. V]. Recent work has shown that if $\widehat{\mathcal{A}}$ is a near-optimal low-rank approximation to $\mathcal{A}$, then $g(\widehat{\mathcal{A}}_{0})$ is a near-optimal low-rank approximation to $g(\mathcal{A})$, for any operator monotone function $g$ satisfying $g(0)=0$; see [32, 33]. Hence, to obtain a low-rank approximation to $\mathcal{A}_{\lambda}$ we compute a Nyström approximation $\widehat{\mathcal{A}}_{0}$ to $\mathcal{A}_{0}$ and then approximate $\mathcal{A}_{\lambda}$ with $g_{\lambda}(\widehat{\mathcal{A}}_{0})$. In our experiments, we set $\lambda=1$ and $d=1$. In this case, the Green’s function is given by
$G(x,y)=\min(x,y)-\frac{xy}{2\pi}.$ (31)
The results are displayed in Fig. 3.
[…]
Figure 3: The contour plots show the exact kernel defined by (31) and the approximated kernels using rank $100$ Nyström approximations. The error plots show the relative error in the Hilbert-Schmidt norm. Optimal denotes the best low-rank approximation error.
5 Conclusions
In this work, we have presented an infinite-dimensional extension of the randomized Nyström approximation. We first established bounds for the finite-dimensional Nyström approximation when the columns of the sketch matrix $\bm{\Omega}$ are drawn independently from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$. Subsequently, through a continuity argument, we provided analogous bounds for the infinite-dimensional extension of the Nyström approximation for Hilbert-Schmidt operators. Additionally, in the process of analyzing the Nyström approximation for Hilbert-Schmidt operators, we have also improved the existing bounds for the randomized SVD for Hilbert-Schmidt operators.
Data availability
All experiments have been performed in MATLAB (version 2020a) on a MacBook Pro with a 2.3 GHz Intel Core i7 processor with 4 cores, and the scripts to reproduce all figures in this paper are publicly available at https://github.com/davpersson/infNystrom.
Appendix A Properties of Gaussian matrices
The following lemma is a consequence of the symmetry of standard normal random variables.
Lemma A.1.
Let $\bm{C}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\bm{D}\in\mathbb{R}^{n_{1}\times n_{2}}$ be two matrices, consider an $m_{2}\times n_{1}$ standard Gaussian matrix $\bm{\Psi}$, and define $\bm{\Phi}=\bm{C}\bm{\bm{\Psi}}\bm{D}$. Then, $\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\bm{0}$.
Proof A.2.
First note that the expectation exists, since each entry of $\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}$ is a linear combination of products of Gaussian random variables, which always has a finite expectation. The distribution of $\bm{\Psi}$ and, in turn, $\bm{\Phi}$ is symmetric, which gives
$-\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\mathbb{E}_{\bm{\Psi}}[-\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\mathbb{E}_{\bm{\Psi}}[(-\bm{\Phi})(-\bm{\Phi})^{*}(-\bm{\Phi})]=\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}],$
implying the result of the lemma.
The following lemma summarizes results on the expected norms of scaled and shifted Gaussian matrices for the Frobenius, Schatten-4, and spectral norms.
Lemma A.3 (Expected norm of shifted Gaussian matrices).
Let $\bm{C}\in\mathbb{R}^{m_{1}\times m_{2}},\bm{D}\in\mathbb{R}^{n_{1}\times n_{2}}$, and $\bm{B}\in\mathbb{R}^{m_{1}\times n_{2}}$ be three arbitrary matrices, and consider an $m_{2}\times n_{1}$ standard Gaussian matrix $\bm{\Psi}$. Then, the following relations hold:
$\mathbb{E}\big{[}\|\bm{B}+\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{\textup{F}}^{2}\big{]}$ $=\|\bm{B}\|_{\textup{F}}^{2}+\|\bm{C}\|_{\textup{F}}^{2}\|\bm{D}\|_{\textup{F}}^{2},$ (32a)
$\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{(4)}^{4}\big{]}$ $=\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{\textup{F}}^{4},$ (32b)
$\mathbb{E}\big{[}\|\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{2}^{2}\big{]}$ $\leq\left(\|\bm{C}\|_{\textup{F}}\|\bm{D}\|_{2}+\|\bm{C}\|_{2}\|\bm{D}\|_{\textup{F}}\right)^{2},$ (32c)
$\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{\textup{F}}^{4}]$ $=2\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{\textup{F}}^{4}.$ (32d)
Proof A.4.
We introduce the matrix $\bm{\Phi}=\bm{C}\bm{\Psi}\bm{D}$ and begin by proving Eq. 32a. Using the linearity of trace and expectation we have:
$\mathbb{E}[\|\bm{B}+\bm{\Phi}\|_{\textup{F}}^{2}]=\|\bm{B}\|_{\textup{F}}^{2}+\mathbb{E}[\|\bm{\Phi}\|_{\textup{F}}^{2}]+2\mathbb{E}[\operatorname{tr}(\bm{B}^{*}\bm{\Phi})]=\|\bm{B}\|_{\textup{F}}^{2}+\|\bm{C}\|_{\textup{F}}^{2}\|\bm{D}\|_{\textup{F}}^{2},$
[…]
$\mathbb{E}[\|\bm{C}\bm{\Psi}\bm{D}\|_{\textup{F}}^{4}]$ $=\mathbb{E}[(\bm{\psi}^{*}\bm{E}^{*}\bm{E}\bm{\psi})^{2}]=\operatorname{Var}(\bm{\psi}^{*}\bm{E}^{*}\bm{E}\bm{\psi})+(\mathbb{E}[\bm{\psi}^{*}\bm{E}^{*}\bm{E}\bm{\psi}])^{2}$
$=2\|\bm{E}^{*}\bm{E}\|_{\textup{F}}^{2}+\operatorname{tr}(\bm{E}^{*}\bm{E})^{2}=2\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{\textup{F}}^{4},$
where we used a standard result that the variance of a quadratic form with Gaussian random vectors is $2\|\bm{E}^{T}\bm{E}\|_{\textup{F}}^{2}$,
Lemma A.5.
Let $\bm{\Omega}_{1}\in\mathbb{R}^{k\times(k+p)}$ be a random matrix whose columns are i.i.d. $\mathcal{N}(0,\widetilde{\bm{K}}_{11})$ and let $\bm{B}\in\mathbb{R}^{k\times n}$ be an arbitrary matrix. Then, the following relation hold for $k\geq 1$ and $p\geq 2$:
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\bm{B}\|_{\textup{F}}^{2}]$ $=\frac{\operatorname{tr}(\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1}\bm{B})}{p-1}=\frac{\|\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1/2}\|_{\textup{F}}^{2}}{p-1},$ (33a)
Additionally, for $p,k\geq 2$ we have:
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]$ $\leq\frac{e^{2}(k+p)}{(p-1)(p+1)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}.$ (34a)
Finally, for $k\geq 1$ and $p\geq 4$ we have
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}]$ $=\frac{(p-1)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}}{p(p-1)(p-3)}\leq k\frac{k+p-1}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2},$ (35a)
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]$ $=\frac{(p-2)\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}+2\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}}{p(p-1)(p-3)}\leq k\frac{kp-2k+2}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}.$ (35b)
Proof A.6.
First, Eq. 33a is proven in [13, Lem. 3.12] and Eq. 34a follows from in [30, Lem. 3.1] and the fact $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]\leq\mathbb{E}[\|\bm{\Psi}^{\dagger}\|_{2}^{2}]\|\bm{K}_{11}^{-1}\|_{2}$ where $\bm{\Psi}$ is a $k\times(k+p)$ standard Gaussian matrix. We proceed to prove the equality in Eq. 35a and introduce the random variable $\bm{X}=\bm{\Omega}_{1}\bm{\Omega}_{1}^{*}\sim\mathcal{W}_{k}(\widetilde{\bm{K}}_{11},k+p)$, such that
$\mathbb{E}\big{[}\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}\big{]}=\mathbb{E}\big{[}\|\bm{X}^{-1}\|_{\textup{F}}^{2}\big{]}=\mathbb{E}\big{[}\operatorname{tr}(\bm{X}^{-2})\big{]}=\frac{\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}}{p(p-3)}+\frac{\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}}{p(p-1)(p-3)},$
where the last equality follows from [23, Thm. 2.4.14]. The equality in Eq. 35b follows from the fact that $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]=\mathbb{E}[\operatorname{tr}(\bm{X}^{-1})^{2}]=\mathbb{E}[\operatorname{tr}\left(\bm{X}^{-1}\otimes\bm{X}^{-1}\right)]$. We then exploit a relation on $\mathbb{E}[\bm{X}^{-1}\otimes\bm{X}^{-1}]$ [23, Thm. 2.4.14] to obtain
$\mathbb{E}\big{[}\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}\big{]}$ $=\frac{(p-2)\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1}\otimes\widetilde{\bm{K}}_{11}^{-1})+\operatorname{tr}(\operatorname{vec}(\widetilde{\bm{K}}_{11}^{-1})\operatorname{vec}(\widetilde{\bm{K}}_{11}^{-1})^{*})+\operatorname{tr}(\bm{C}_{k\times k}(\widetilde{\bm{K}}_{11}^{-1}\otimes\widetilde{\bm{K}}_{11}^{-1}))}{p(p-1)(p-3)}$
$=\frac{(p-2)\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}+2\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}}{p(p-1)(p-3)},$
where the second equality comes from the relation $\operatorname{tr}(\bm{C}_{k\times k}(\widetilde{\bm{K}}_{11}^{-1}\otimes\widetilde{\bm{K}}_{11}^{-1}))=\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}$, where $\bm{C}_{k\times k}$ is the commutation matrix. [^1] The commutation matrix $\bm{C}_{k\times k}\in\mathbb{R}^{k^{2}\times k^{2}}$ is a $k\times k$ block matrix, with blocks of size $k\times k$. The $(i,j)$-block of $\bm{C}_{k\times k}$ is the matrix $\bm{E}_{ji}$ with entries $(\bm{E}_{ji})_{k\ell}=\delta_{jk}\delta_{i\ell}$. Finally, the inequalities in Eq. 35a and Eq. 35b follow from standard norm inequalities.
The next lemma is a generalization and a consequence of [20, Prop. 10.4], which provides tailbounds on the Frobenius and spectral norms of pseudoinverted standard Gaussian matrices.
Lemma A.7 (Norm bounds for a pseudoinverted scaled Gaussian matrix).
Let $\bm{\Omega}_{1}\in\mathbb{R}^{k\times(k+p)}$ be a Gaussian matrix whose columns are i.i.d. and follow a multivariate Gaussian distribution with mean zero and covariance matrix $\widetilde{\bm{K}}_{11}$. Then, the following relations hold for $p\geq 4$ and all $t\geq 1$:
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}>\sqrt{\frac{3\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p+1}}t\right\}$ $\leq t^{-p},$ (36a)
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{2}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}}{p+1}t\right\}$ $\leq t^{-(p+1)},$ (36b)
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}}}{p+1}t\right\}$ $\leq t^{-(p+1)}.$ (36c)
Proof A.8.
Note that Eq. 36a is a restatement of [8, Lem. 3]. Because $\bm{\Omega}_{1}=\widetilde{\bm{K}}_{11}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix, it follows that $\|\bm{\Psi}^{\dagger}\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}\leq\|\bm{\Psi}^{\dagger}\|_{2}\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}$ for any Schatten-$s$ norm. Moreover, the combination with [20, Prop. 10.4] implies
$\mathbb{P}\left\{\|\bm{\Psi}^{\dagger}\|_{2}>\frac{e\sqrt{(k+p)}}{p+1}t\right\}\leq t^{-(p+1)},$
which yield the bounds Eqs. 36b and 36c using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ and $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{(4)}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}$, respectively.